The sidecap package (two column layout)

In this document, we used the following options of the sidecap package:

\usepackage
[outercaption,raggedright]
{sidecap}

Note that fig. 2 is an SCfigure* environment. The sidecap package inherits the standard L^{A}TEX “feature” that two-column floats and one-column floats may be out of order.

[figure]

Fig. 1: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure] Fig. 3: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure] Fig. 4: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure] Tab. 1: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Fig. 2: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]

Tab. 2: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

1 2 3 4 5 6
2 4 6 8 10 12
3 6 9 12 15 18
4 8 12 16 20 24
5 10 15 20 25 30
6 12 18 24 30 36
7 14 21 28 35 42
8 16 24 32 40 48
9 18 27 36 45 54
10 20 30 40 50 60
11 22 33 44 55 66
12 24 36 48 60 72
13 26 39 52 65 78
14 28 42 56 70 84

Tab. 3: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!  1 2 3 4 5 6 2 4 6 8 10 12 3 6 9 12 15 18 4 8 12 16 20 24 5 10 15 20 25 30 6 12 18 24 30 36 7 14 21 28 35 42 8 16 24 32 40 48 9 18 27 36 45 54 10 20 30 40 50 60 11 22 33 44 55 66 12 24 36 48 60 72 13 26 39 52 65 78 14 28 42 56 70 84

[figure] Fig. 5: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Fig. 6: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

[figure]